OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

for part 1 you can use elimination method. simply multiply the first equation by 2 then add it to the second equation

OpenStudy (anonymous):

part 2 would be subjective

OpenStudy (austinl):

I would just use substitution for this one. It is very straightforward.

OpenStudy (austinl):

We have \[x-y=13\] \[2x+y=2\] I would isolate the y in the second equation, \[y=-2x+2\] and then plug that into the first equation, \[x-(-2x+2)=13\]

OpenStudy (austinl):

Then you solve for x, \[x=11\] Then you take that result and plug it back into one of the equations and solve for y.

OpenStudy (austinl):

Does that make sense?

OpenStudy (anonymous):

yeah

OpenStudy (austinl):

So what did you get for your y value?

OpenStudy (anonymous):

24 not sure if its right if not show me how to plug it into y

OpenStudy (austinl):

We got x=11 correct? So we would put that back into an equation to solve for y. x-y=13 (11)-y=13 -y=2 y=-2

OpenStudy (anonymous):

okay thax

OpenStudy (austinl):

I don't want to be that guy, but it is considered polite to award a medal to people who give good answers and guide you to the correct answer.

OpenStudy (anonymous):